OpenStudy (anonymous):

H(x) = x^3 + 4x^2 -4x-16 I have found all of the intercepts but I dont know which equation to plug numbers into in order to graph this function, do I plug into the original function listed above, or my final simplified answer H(x)= (x+4)(x+2)(x-2). Please help thanks!

OpenStudy (anonymous):

H(x)= (x+4)(x+2)(x-2) = x^3 + 4x^2 -4x-16 so it makes no difference.

OpenStudy (anonymous):

then why are my graphs turning out ridiculous? could you draw what this graph should look like?

OpenStudy (anonymous):

woah thats awesome thank you!

OpenStudy (anonymous):

ur welcome